OpenStudy (anonymous):

If (7^2^)^p = 7^6, what is the value of p? 2 3 4 8

OpenStudy (chillhill):

3

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \left(7^2\right)^p = 7^6\] \[\Large 7^{2*p} = 7^6\]

OpenStudy (anonymous):

3?